find the equation of the line perpendicular to y=-7/5x+100, passing through the points (-14,-52)
Answer by CharlesG2(834)

find the equation of the line perpendicular to y=-7/5x+100, passing through the points (-14,-52)
this line has slope m of -7/5
the perpendicular line needs to be -1/m and the product of the slopes needs to be -1
-1/m = -1/(-7/5) = 5/7 which is the slope of the perpendicular line
and -7/5 * 5/7 = -1
now need to solve for new b (y-intercept or vertical-intercept or point (0,b))
5/7 * -14 + b = -52
-70/7 + b = -52
-10 + b = -52
b = -42
equation of line is y = 5/7 * x - 42